# SpiderCat: A Catalog of Compact Binary Millisecond Pulsars

### v1

@@ 3.5.2 Optical/IR
PSR J1705-1903 $5.8^{+2.9}_{-2.4}$ $3.6\pm 0.8$ $14.5^{+25.3}_{-10.5}$ 19.02 $5.2^{+1.1}_{-0.9}$ 0.30
PSR J1720-0533 $1.3^{+0.5}_{-0.5}$
PSR J1723-2837 $0.7^{+0.3}_{-0.3}$ $229.3\pm 2.8$ $14.2^{+14.0}_{-14.2}$ 12.34 $3.1^{+1.1}_{-0.7}$0.31
PSR J1731-1847 $4.8^{+1.9}_{-1.9}$ $1.6\pm 0.4$ $4.5^{+6.4}_{-4.5}$ $5.2\pm 1.1$ $14.1^{+19.3}_{-10.1}$ 18.50 $5.1^{+1.1}_{-0.7}$ 0.65
PSR J1745-23 $7.9^{+3.2}_{-3.2}$ 16.65 $2.1^{+1.1}_{-0.7}$ 0.78
PSR J1745+1017 $1.2^{+0.5}_{-0.5}$ $7.6\pm 0.6$ $1.3^{+1.5}_{-0.9}$
PSR J1803-6707 $1.4^{+0.6}_{-0.6}$ $5.9\pm 2.0$ $1.5^{+2.4}_{-1.5}$ $4.8\pm 0.5$ $1.2^{+1.4}_{-0.8}$ 20.02 $9.2^{+1.1}_{-0.7}$ 0.53
PSR J1805+0615 $3.9^{+1.6}_{-1.6}$ $5.3\pm 0.5$ $9.5^{+11.1}_{-6.4}$0.76
PSR J1810+1744 $2.4^{+0.9}_{-0.9}$ $1.8\pm 0.3$ $1.2^{+1.6}_{-1.2}$ $23.2\pm 0.9$ $15.5^{+16.1}_{-10.1}$ 19.68 $7.8^{+1.1}_{-0.7}$0.25
4FGL J1813.5+2819 $3.2^{+1.7}_{-0.9}$ $2.3\pm 0.3$ $2.9^{+4.9}_{-1.6}$ 18.79 $6.3^{+0.7}_{-0.9}$ 0.29
PSR J1814+0045g
PSR J1816+4510 $4.3^{+1.7}_{-1.7}$ $0.2\pm 0.1$ $0.5^{+0.8}_{-0.5}$ $10.6\pm 0.5$ $23.8^{+25.0}_{-15.6}$ 17.95 $4.8^{+1.1}_{-0.7}$0.33
4FGL J1819.4-1102 $4.5^{+3.3}_{-2.4}$ $8.2\pm 2.8$ $19.6^{+59.5}_{-16.7}$ 12.28 $-1.0^{+1.6}_{-1.2}$0.24
4FGL J1824.2+1231 $2.4^{+2.0}_{-0.9}$ $3.3\pm 0.9$ $2.3^{+7.3}_{-1.6}$ 18.90 $7.0^{+1.0}_{-1.3}$ 0.36
PSR J1830-0106g -5.87
PSR J1833-3840 $4.7^{+1.9}_{-1.9}$ $2.8\pm 0.4$ $7.3^{+9.3}_{-5.1}$
4FGL J1838.2+3223 $2.2^{+2.4}_{-1.2}$ $2.7\pm 0.6$ $1.6^{+6.9}_{-1.3}$ 20.45 $8.7^{+1.8}_{-1.6}$ 0.19
PSR J1838+1507g
PSR J1847+0342g
PSR J1849+0304g
4FGL J1853.6-0620 $5.4^{+2.8}_{-2.3}$ $4.9\pm 1.4$ $17.0^{+33.5}_{-13.0}$1.97
4FGL J1859.2-0706 $5.8^{+2.6}_{-2.1}$ $4.8\pm 1.1$ $19.4^{+31.5}_{-13.3}$ 18.46 $4.6^{+1.0}_{-0.8}$ 0.01
PSR J1859+0313g $3.1^{+1.2}_{-1.2}$
4FGL J1901.8-0718 $3.5^{+3.1}_{-2.1}$ $3.7\pm 0.7$ $5.3^{+16.9}_{-4.6}$ 19.48 $6.8^{+2.0}_{-1.4}$ 0.31
PSR J1908+2105 $2.6^{+1.0}_{-1.0}$ $2.4\pm 0.6$ $2.0^{+2.9}_{-2.0}$ $4.9\pm 0.8$ $4.0^{+5.0}_{-2.8}$ 19.40 $7.3^{+1.1}_{-0.7}$ 0.38
PSR J1910-5320 $1.0^{+0.4}_{-0.4}$ $19.1\pm 2.1$ $2.3^{+2.7}_{-2.3}$ $3.1\pm 0.5$ $0.4^{+0.5}_{-0.3}$ 18.79 $8.8^{+1.1}_{-0.7}$ 0.25
PSR J1919+0126g
PSR J1919+1502g $4.0^{+3.0}_{-1.9}$
PSR J1928+1245 $6.1^{+2.4}_{-2.4}$ 14.64 $0.7^{+1.1}_{-0.7}$0.30
PSR J1931+1428g
PSR J1932+2121 $4.9^{+2.9}_{-2.0}$ 8.70 -2.36
PSR J1946-5403 $1.1^{+0.5}_{-0.5}$ $1.0\pm 0.3$ $0.2^{+0.2}_{-0.2}$ $9.8\pm 0.5$ $1.6^{+1.6}_{-1.0}$
PSR J1947-1120 $5.7^{+2.0}_{-1.3}$ $3.3\pm 0.6$ $12.9^{+15.1}_{-6.7}$ 16.59 $2.8^{+0.6}_{-0.7}$ 0.72
PSR J1953+1006g
PSR J1957+2516 $2.7^{+1.1}_{-1.1}$ $4.0\pm 1.3$ $3.4^{+5.5}_{-2.6}$ 12.49 $0.4^{+1.1}_{-0.7}$0.82
PSR B1957+20 $1.7^{+0.7}_{-0.7}$ $6.8\pm 0.3$ $2.4^{+2.5}_{-2.4}$ $15.7\pm 0.9$ $5.6^{+6.0}_{-3.7}$ 19.53 $8.3^{+1.1}_{-0.7}$ 0.30
PSR J2003+3032g
PSR J2017-1614 $1.4^{+0.6}_{-0.6}$ $1.8\pm 0.3$ $0.4^{+0.6}_{-0.4}$ $6.5\pm 0.6$ $1.6^{+1.8}_{-1.1}$0.23
PSR J2039-5617 $1.7^{+0.7}_{-0.7}$ $14.2\pm 0.8$ $4.9^{+5.3}_{-4.9}$ $15.4\pm 0.6$ $5.4^{+5.6}_{-3.5}$ 18.74 $7.6^{+1.1}_{-0.7}$ 0.41
PSR J2047+1053 $2.8^{+1.1}_{-1.1}$ $0.7\pm 0.2$ $0.6^{+0.9}_{-0.6}$ $4.3\pm 0.6$ $4.0^{+4.9}_{-2.7}$ 22.15 $9.9^{+1.1}_{-0.7}$ 0.34
PSR J2051-0827 $1.5^{+0.6}_{-0.6}$ $0.8\pm 0.3$ $0.2^{+0.3}_{-0.2}$ $2.5\pm 0.3$ $0.6^{+0.8}_{-0.4}$
PSR J2052+1219 $3.9^{+1.6}_{-1.6}$ $4.6\pm 0.6$ $8.4^{+10.1}_{-5.7}$ 21.54 $8.6^{+1.1}_{-0.7}$0.55
4FGL J2054.2+6904 $2.6^{+1.8}_{-1.0}$ $11.1\pm 2.2$ $8.9^{+22.3}_{-8.9}$ $4.2\pm 0.5$ $3.4^{+7.7}_{-2.3}$ 19.38 $7.3^{+1.1}_{-1.2}$ 0.29
PSR J2055+3829 $4.6^{+1.8}_{-1.8}$2.21
PSR J2055+1545 $3.7^{+1.5}_{-1.5}$ $2.3\pm 0.6$ $3.7^{+5.4}_{-2.7}$ 20.74 $7.9^{+1.1}_{-0.7}$ 0.54
PSR J2115+5448 $3.1^{+1.2}_{-1.2}$ $1.3\pm 0.3$ $1.5^{+2.2}_{-1.5}$ $7.0\pm 0.7$ $8.1^{+9.3}_{-5.5}$
PSR J2129-0429 $1.4^{+0.6}_{-0.6}$ $23.2\pm 0.6$ $5.3^{+5.4}_{-5.3}$ $6.8\pm 0.5$ $1.6^{+1.7}_{-1.0}$ 17.34 $6.6^{+1.1}_{-0.7}$ 0.71
PSR J2214+3000 $1.7^{+0.7}_{-0.7}$ $4.2\pm 0.3$ $1.4^{+1.6}_{-1.4}$ $32.6\pm 0.7$ $10.9^{+11.0}_{-7.1}$
PSR J2215+5135 $2.8^{+1.1}_{-1.1}$ $14.4\pm 0.8$ $13.3^{+14.2}_{-13.3}$ $18.0\pm 0.8$ $16.6^{+17.4}_{-10.9}$ 19.30 $7.1^{+1.1}_{-0.7}$ 0.40
3FGL J2221.6+6507 $2.3^{+0.6}_{-0.4}$ 15.85 $4.1^{+0.4}_{-0.5}$ 0.21
PSR J2234+0944 $1.6^{+0.6}_{-0.6}$ $10.0\pm 0.6$ $3.0^{+3.3}_{-2.0}$
PSR J2241-5236 $1.0^{+0.4}_{-0.4}$ $4.9\pm 0.3$ $0.5^{+0.6}_{-0.5}$ $25.0\pm 1.1$ $2.8^{+2.9}_{-1.8}$
PSR J2256-1024 $1.3^{+0.5}_{-0.5}$ $4.0\pm 0.4$ $0.9^{+1.0}_{-0.9}$ $8.2\pm 0.5$ $1.7^{+1.9}_{-1.1}$ 21.75 $11.1^{+1.1}_{-0.7}$ 0.54
PSR J2333-5526 $2.5^{+1.0}_{-1.0}$ $7.2\pm 1.4$ $5.3^{+7.0}_{-5.3}$ $4.3\pm 0.4$ $3.1^{+3.6}_{-2.1}$ 21.18 $9.2^{+1.1}_{-0.7}$ 1.06
PSR J2339-0533 $0.8^{+0.3}_{-0.3}$ $26.0\pm 0.8$ $1.8^{+1.8}_{-1.8}$ $29.2\pm 0.8$ $2.0^{+2.0}_{-1.3}$ 18.15 $8.8^{+1.1}_{-0.7}$1.30
@@ 3.6 Spin-orbit
Figure 9: Spider spin periods as a function of orbital period (panel a: RBs, panel b: BWs). Data points are color-coded by spin-down luminosity (black indicates no available estimate). Most spiders cluster along two axes or groups (labeled 1 and 2, and highlighted by magenta and green dashed boxes, respectively): Group 1 features fast spins and longer orbital periods, while Group 2 comprises systems with short orbital periods and a broader range of spin periods. Upper panels display box plots of selected parameters for the two groups: spin-down luminosity (left), characteristic age (middle), and magnetic field strength at the light cylinder (right).
@@
Spiders appear to cluster along two axes in the spin period versus orbital period plane (Fig.  9). The majority of sources fall into one of two groups: either fast-spinning pulsars (1.5–2.5 ms) with longer orbital periods (Group 1 in Fig. 9), or systems with short orbital periods ($\sim$3 hr for BWs, $\sim$5 hr for RBs) and a broader range of spin periods (1.5–5 ms; Group 2 in Fig. 9). These groups also differ significantly in spin-down luminosity ($\sigma\sim 3-3.5$) [^9] Medians $\bar{\dot{E}}_{\rm RB,1}=10^{34.9}$ erg/s and $\bar{\dot{E}}_{\rm RB,2}=10^{34.4}$ erg/s (Mann-Whitney $U=76$, $n_{1}=7$, $n_{2}=12$, $p<0.0026$, two-tailed) for RBs; $\bar{\dot{E}}_{\rm BW,1}=10^{34.9}$ erg/s and $\bar{\dot{E}}_{\rm BW,2}=10^{34.1}$ erg/s (Mann-Whitney $U=212$, $n_{1}=12$, $n_{2}=20$, $p<0.0004$, two-tailed) for BWs., characteristic age ($\sigma\sim 1.6-2.7$) [^10] Medians $\bar{\tau}_{\rm RB,1}=10^{9.35}$ yr and $\bar{\tau}_{\rm RB,2}=10^{9.53}$ yr (Mann-Whitney $U=22$, $n_{1}=7$, $n_{2}=12$, $p<0.1$, two-tailed) for RBs; $\bar{\tau}_{\rm BW,1}=10^{9.33}$ yr and $\bar{\tau}_{\rm BW,2}=10^{9.76}$ yr (Mann-Whitney $U=49$, $n_{1}=12$, $n_{2}=20$, $p<0.006$, two-tailed) for BWs., and magnetic field strength at the light cylinder ($\sigma\sim 2.8-3.7$) [^11] Medians $\bar{B}_{\rm LC,RB,1}=10^{5.28}$ G and $\bar{B}_{\rm LC,RB,2}=10^{4.94}$ G (Mann-Whitney $U=74$, $n_{1}=7$, $n_{2}=12$, $p<0.005$, two-tailed) for RBs; $\bar{B}_{\rm LC,BW,1}=10^{5.32}$ G and $\bar{B}_{\rm LC,BW,2}=10^{4.82}$ G (Mann-Whitney $U=215$, $n_{1}=12$, $n_{2}=20$, $p<0.0002$, two-tailed) for BWs., as shown in the upper panels of Fig. 9. Here, we defined the magnetic field strength at the light cylinder as $B_{\rm LC}=B(R/R_{\rm LC})^{3}$, where $B$ is the surface magnetic field strength defined in Section 3.2, $R=10$ km, and $R_{\rm LC}=cP/2\pi$. Compared to Group 2, Group 1 exhibits higher spin-down luminosities, younger characteristic ages, and stronger magnetic fields at the light cylinder. These trends are consistent across both RBs (Fig. 9, panel a) and BWs (Fig. 9, panel b).
@@
According to binary evolution models of spider systems, slower spin periods may result from lower accretion efficiency during the recycling phase or from a higher initial neutron star mass (e.g., T. M. Tauris et al., 2012; H.-L. Chen et al., 2013; A. Kar et al., 2024; D. Misra et al., 2025a, and references therein). These long-$P$ systems would also exhibit lower spin-down luminosities ($\dot{E}\propto P^{-3}$), leading to reduced companion irradiation. Weaker irradiation inhibits mass loss from the binary, thereby limiting orbital widening. As a result, Group 2 spiders would maintain a narrow range of orbital periods while showing a wider distribution of spin periods, shaped by variations in initial neutron star mass or accretion efficiency. In contrast, spiders in Group 1 likely underwent more efficient accretion during recycling phase, reaching spin periods of $\sim$2 ms. These systems display higher spin-down luminosities and likely stronger companion irradiation, which leads to wider orbits and a broader range of orbital periods depending on irradiation efficiency.
A few outliers exist outside these two main groups or clusters. Among RBs, the systems not shown in Fig. 9 (panel a) include huntsmans and two long spin period sources: PSR J1932$+$2121 and PSR J2129--0429, which are mildly recycled pulsars. The three sources shown in the figure outside the two groups are PSR J1302--3258, PSR J1242--4712 and PSR J1816$+$4510.
PSR J1302--3258 has $P_{\rm b}=18.8$ h, $P=3.77$ ms, $M_{\rm c,min}=0.15M_{\odot}$ and shows eclipses at low radio frequencies, fulfilling our RB definition (Sec. 1.2). However, the eclipses last for only $\sim$10of the orbit, and the system lacks an optical counterpart despite its apparent proximity ($d\approx 1.4$ kpc from DM and YMW16, Sec. 3.4), raising doubts about its RB identification (P. Bangale et al., 2024). Similarly, PSR J1242--4712, with $P_{\rm b}=7.8$ h, $P=5.31$ ms, $M_{\rm c,min}=0.09M_{\odot}$, shows eclipses only a short fraction of its orbit ($\sim$2%) and lacks an optical counterpart. However, its greater distance ($d\approx 2.5$ kpc from DM and YMW16, Sec. 3.4) could still accommodate a relatively cool, undetected companion (Tsimilar-$\sim$4000 K, A. Ghosh et al., 2024). Thus, deeper optical-IR observations are needed to establish the nature of the companions to PSR J1302--3258 and PSR J1242--4712. Optical spectroscopy of PSR J1816$+$4510 indicates that its companion is a proto-white dwarf (D. L. Kaplan et al., 2013). Therefore, all RB outliers in the $P-P_{\rm b}$ diagram appear to be special cases, with their atypical companions or unusually long spin periods implying peculiar evolutionary paths.

### v4

@@ 3.5.2 Optical and Infrared
PSR J1705-1903 $\sim$2.3 YMW16 $3.6\pm 0.8$ $2.4^{+3.3}_{-1.7}$ 19.02 $7.2^{+1.1}_{-0.7}$ 0.30
PSR J1720-0533 $\sim$1.3 NE2001
PSR J1723-2837 $0.91^{+0.04}_{-0.03}$ Gaia $229.3\pm 2.8$ XMM $22.5^{+2.2}_{-2.0}$ 12.34 $2.55^{+0.09}_{-0.09}$ -0.31
PSR J1731-1847 $\sim$4.8 YMW16 $1.6\pm 0.4$ XMM $4.5^{+6.4}_{-3.3}$ $5.2\pm 1.1$ $14.1^{+19.3}_{-10.1}$ 18.50 $5.1^{+1.1}_{-0.7}$ 0.65
PSR J1745-23 $\sim$7.9 YMW16 16.65 $2.1^{+1.1}_{-0.7}$ 0.78
PSR J1745+1017 $\sim$1.2 YMW16 $7.6\pm 0.6$ $1.3^{+1.5}_{-0.9}$
PSR J1803-6707 $\sim$1.4 YMW16 $5.9\pm 2.0$ eRosita $1.5^{+2.4}_{-1.1}$ $4.8\pm 0.5$ $1.2^{+1.4}_{-0.8}$ 20.02 $9.2^{+1.1}_{-0.7}$ 0.53
PSR J1805+0615 $\sim$3.9 YMW16 $5.3\pm 0.5$ $9.5^{+11.1}_{-6.4}$ -0.76
PSR J1810+1744 $\sim$2.4 YMW16 $1.8\pm 0.3$ Chandra $1.2^{+1.6}_{-0.8}$ $23.2\pm 0.9$ $15.5^{+16.1}_{-10.1}$ 19.68 $7.8^{+1.1}_{-0.7}$ -0.25
4FGL J1813.5+2819 $3.2^{+1.7}_{-0.9}$ Gaia $2.3\pm 0.3$ $2.9^{+4.9}_{-1.6}$ 18.79 $6.3^{+0.7}_{-0.9}$ 0.29
PSR J1814+0045g $\sim$8.9 YMW16
PSR J1816+4510 $\sim$4.3 YMW16 $0.2\pm 0.1$ Chandra $0.5^{+0.8}_{-0.4}$ $10.6\pm 0.5$ $23.8^{+25.0}_{-15.6}$ 17.95 $4.8^{+1.1}_{-0.7}$ -0.33
4FGL J1819.4-1102 $4.5^{+3.3}_{-2.4}$ Gaia $8.2\pm 2.8$ $19.7^{+59.2}_{-16.8}$ 12.28 $1.0^{+1.6}_{-1.2} -$0.24
4FGL J1824.2+1231 $2.4^{+2.0}_{-0.9}$ Gaia $3.3\pm 0.9$ $2.2^{+7.2}_{-1.6}$ 18.90 $7.0^{+1.0}_{-1.3}$ 0.36
PSR J1830-0106g $\sim$5.0 YMW16 -5.87
PSR J1833-3840 $\sim$4.7 YMW16 $2.8\pm 0.4$ $7.3^{+9.3}_{-5.1}$
4FGL J1838.2+3223 $2.2^{+2.5}_{-1.2}$ Gaia $2.7\pm 0.6$ $1.5^{+6.9}_{-1.3}$ 20.45 $8.7^{+1.8}_{-1.6}$ 0.19
PSR J1838+1507g $\sim$2.4 YMW16
PSR J1847+0342g $\sim$2.9 YMW16
PSR J1849+0304g $\sim$4.5 YMW16
4FGL J1853.6-0620 $5.4^{+2.8}_{-2.3}$ Gaia $4.9\pm 1.4$ $17.0^{+33.7}_{-13.0}$ -1.97
4FGL J1859.2-0706 $5.8^{+2.6}_{-2.1}$ Gaia $4.8\pm 1.1$ $19.4^{+31.4}_{-13.4}$ 18.46 $4.6^{+1.0}_{-0.8}$ 0.01
PSR J1859+0313g $\sim$3.1 YMW16
4FGL J1901.8-0718 $3.4^{+3.1}_{-2.1}$ Gaia $3.7\pm 0.7$ $5.2^{+17.0}_{-4.5}$ 19.48 $6.8^{+2.0}_{-1.4}$ 0.31
PSR J1908+2105 $\sim$2.6 YMW16 $2.4\pm 0.6$ XMM $2.0^{+2.9}_{-1.4}$ $4.9\pm 0.8$ $4.0^{+5.0}_{-2.8}$ 19.40 $7.3^{+1.1}_{-0.7}$ 0.38
PSR J1910-5320 $\sim$1.0 YMW16 $19.1\pm 2.1$ Chandra $2.3^{+2.7}_{-1.5}$ $3.1\pm 0.5$ $0.4^{+0.5}_{-0.3}$ 18.79 $8.8^{+1.1}_{-0.7}$ 0.25
PSR J1919+0126g $\sim$6.3 YMW16
PSR J1919+1502g $\sim$6.0 YMW16
PSR J1928+1245 $\sim$6.1 YMW16 14.64 $0.7^{+1.1}_{-0.7}$ -0.30
PSR J1931+1428g $\sim$6.2 YMW16
PSR J1932+2121 $\sim$5.1 YMW16 8.70 -2.36
PSR J1946-5403 $\sim$1.1 YMW16 $1.0\pm 0.3$ XMM $0.2^{+0.2}_{-0.1}$ $9.8\pm 0.5$ $1.6^{+1.6}_{-1.0}$
PSR J1947-1120 $3.3\pm 0.6$ 16.59 0.72
PSR J1953+1006g $\sim$2.5 YMW16
PSR J1957+2516 $\sim$2.7 YMW16 $4.0\pm 1.3$ $3.4^{+5.5}_{-2.6}$ 12.49 $0.4^{+1.1}_{-0.7}$ -0.82
PSR B1957+20 $\sim$1.7 YMW16 $6.8\pm 0.3$ XMM $2.4^{+2.5}_{-1.6}$ $15.7\pm 0.9$ $5.6^{+6.0}_{-3.7}$ 19.53 $8.3^{+1.1}_{-0.7}$ 0.30
PSR J2003+3032g $\sim$6.8 YMW16
PSR J2017-1614 $\sim$1.4 YMW16 $1.8\pm 0.3$ XMM $0.4^{+0.6}_{-0.3}$ $6.5\pm 0.6$ $1.6^{+1.8}_{-1.1}$ -0.23
PSR J2039-5617 $\sim$1.7 YMW16 $14.2\pm 0.8$ XMM $4.9^{+5.3}_{-3.3}$ $15.4\pm 0.6$ $5.4^{+5.6}_{-3.5}$ 18.74 $7.6^{+1.1}_{-0.7}$ 0.41
PSR J2047+1053 $\sim$2.8 YMW16 $0.7\pm 0.2$ Chandra $0.6^{+0.9}_{-0.4}$ $4.3\pm 0.6$ $4.0^{+4.9}_{-2.7}$ 22.15 $9.9^{+1.1}_{-0.7}$ 0.34
PSR J2051-0827 $\sim$1.5 YMW16 $0.8\pm 0.3$ XMM $0.2^{+0.3}_{-0.2}$ $2.5\pm 0.3$ $0.6^{+0.8}_{-0.4}$
PSR J2052+1219 $\sim$3.9 YMW16 $4.6\pm 0.6$ $8.4^{+10.1}_{-5.7}$ 21.54 $8.6^{+1.1}_{-0.7}$ -0.55
4FGL J2054.2+6904 $2.6^{+1.9}_{-1.0}$ Gaia $11.1\pm 2.2$ Swift $8.9^{+22.7}_{-6.3}$ $4.2\pm 0.5$ $3.3^{+7.8}_{-2.3}$ 19.38 $7.3^{+1.1}_{-1.2}$ 0.29
PSR J2055+3829 $\sim$4.6 YMW16 -2.21
PSR J2055+1545 $\sim$3.7 YMW16 $2.3\pm 0.6$ $3.7^{+5.4}_{-2.7}$ 20.74 $7.9^{+1.1}_{-0.7}$ 0.54
PSR J2115+5448 $\sim$3.1 YMW16 $1.3\pm 0.3$ XMM $1.5^{+2.2}_{-1.1}$ $7.0\pm 0.7$ $8.1^{+9.3}_{-5.5}$
PSR J2129-0429 $2.0^{+0.3}_{-0.2}$ Gaia $23.2\pm 0.6$ XMM $10.8^{+4.0}_{-2.6}$ $6.8\pm 0.5$ $3.2^{+1.4}_{-0.9}$ 17.34 $5.9^{+0.3}_{-0.3}$ 0.71
PSR J2214+3000 $\sim$1.7 YMW16 $4.2\pm 0.3$ XMM $1.4^{+1.6}_{-0.9}$ $32.6\pm 0.7$ $10.9^{+11.0}_{-7.1}$
PSR J2215+5135 $\sim$2.8 YMW16 $14.4\pm 0.8$ XMM $13.3^{+14.2}_{-8.7}$ $18.0\pm 0.8$ $16.6^{+17.4}_{-10.9}$ 19.30 $7.1^{+1.1}_{-0.7}$ 0.40
3FGL J2221.6+6507 $2.3^{+0.6}_{-0.4}$ Gaia 15.85 $4.1^{+0.4}_{-0.5}$ 0.21
PSR J2234+0944 $\sim$1.6 YMW16 $10.0\pm 0.6$ $3.0^{+3.3}_{-2.0}$
PSR J2241-5236 $\sim$1.0 YMW16 $4.9\pm 0.3$ XMM $0.5^{+0.6}_{-0.4}$ $25.0\pm 1.1$ $2.8^{+2.9}_{-1.8}$
PSR J2256-1024 $\sim$1.3 YMW16 $4.0\pm 0.4$ Chandra $0.9^{+1.0}_{-0.6}$ $8.2\pm 0.5$ $1.7^{+1.9}_{-1.1}$ 21.75 $11.1^{+1.1}_{-0.7}$ 0.54
PSR J2333-5526 $\sim$2.5 YMW16 $7.2\pm 1.4$ XMM $5.3^{+7.0}_{-3.7}$ $4.3\pm 0.4$ $3.1^{+3.6}_{-2.1}$ 21.18 $9.2^{+1.1}_{-0.7}$ 1.06
PSR J2339-0533 $\sim$0.8 YMW16 $26.0\pm 0.8$ XMM $1.8^{+1.8}_{-1.1}$ $29.2\pm 0.8$ $2.0^{+2.0}_{-1.3}$ 18.15 $8.8^{+1.1}_{-0.7}$ -1.30
@@ 3.6 Spin-orbit
Figure 9: Spider spin periods as a function of orbital period (panel (a): RBs, panel (b): BWs). Data points are color coded by K-means clusters (black indicates no available estimate for $\dot{P}$): Cluster 1 (violet data points) features fast spins and a wider range in orbital periods, while Cluster 2 (green data points) comprises systems with short orbital periods and a broader range of spin periods. Upper panels display box plots of derived parameters for the two clusters: spin-down luminosity (left), characteristic age (middle), and magnetic field strength at the light cylinder (right).
Spiders appear to cluster along two axes in the spin period versus orbital period plane (Figure 9). The majority of sources fall into one of two groups: either fast-spinning pulsars (<<3 ms) with a wide range of orbital periods, or systems with short orbital periods ($\sim$3 hr for BWs, $\sim$5 hr for RBs) and a broader range of spin periods (2–5 ms).
To quantify this, we performed a K-means clustering on a spider-type separated dataset including parameters $P$, $\dot{P}$, $P_{\rm b}$, and the projected semi-major axis of the orbit ($A_{1}$) acquired from ATNF. From this analysis, we exclude HMs and two long spin period sources: PSR J1932++2121 and PSR J2129-0429, which are mildly recycled pulsars. The results for two clusters are shown in Figure 9 for both RBs and BWs, possibly indicating differing source populations. We also derived distributions for the spin-down luminosity, characteristic age, and the magnetic strength at the light cylinder for both clusters (top panels in Figure 9). Here, we defined the magnetic field strength at the light cylinder as $B_{\rm LC}=B(R/R_{\rm LC})^{3}$, where $B$ is the surface magnetic field strength defined in Section 3.2, $R=10$ km, and $R_{\rm LC}=cP/2\pi$.
@@
These clusters also differ significantly in spin-down luminosity ($\sigma\sim 2.6-3.1$), [^11] Medians $\bar{\dot{E}}_{\rm RB,1}=10^{34.8}$ erg s-1 and $\bar{\dot{E}}_{\rm RB,2}=10^{34.2}$ erg s-1 (Mann-Whitney $U=101$, $n_{1}=11$, $n_{2}=11$, $p<0.009$, two-tailed) for RBs; $\bar{\dot{E}}_{\rm BW,1}=10^{34.8}$ erg s-1 and $\bar{\dot{E}}_{\rm BW,2}=10^{34.2}$ erg s-1 (Mann-Whitney $U=265$, $n_{1}=13$, $n_{2}=25$, $p<0.002$, two-tailed) for BWs. and magnetic field strength at the light cylinder ($\sigma\sim 3.1-3.5$), [^12] Medians $\bar{B}_{\rm LC,RB,1}=10^{5.3}$ G and $\bar{B}_{\rm LC,RB,2}=10^{4.7}$ G (Mann-Whitney $U=109$, $n_{1}=11$, $n_{2}=11$, $p<0.002$, two-tailed) for RBs; $\bar{B}_{\rm LC,BW,1}=10^{5.2}$ G and $\bar{B}_{\rm LC,BW,2}=10^{4.8}$ G (Mann-Whitney $U=277$, $n_{1}=13$, $n_{2}=25$, $p<0.0005$, two-tailed) for BWs. but not in characteristic age ($\sigma\sim 0.3-1.5$), [^13] Medians $\bar{\tau}_{\rm RB,1}=10^{9.5}$ yr and $\bar{\tau}_{\rm RB,2}=10^{9.5}$ yr (Mann-Whitney $U=56$, $n_{1}=11$, $n_{2}=11$, $p<0.8$, two-tailed) for RBs; $\bar{\tau}_{\rm BW,1}=10^{9.5}$ yr and $\bar{\tau}_{\rm BW,2}=10^{9.7}$ yr (Mann-Whitney $U=113$, $n_{1}=13$, $n_{2}=25$, $p<0.13$, two-tailed) for BWs. as shown in the upper panels of Figure 9. Overall, compared to Cluster 2, Cluster 1 exhibits higher spin-down luminosities and stronger magnetic fields at the light cylinder. These trends are consistent across both BWs (Figure 9, panel (a)) and RBs (Figure 9, panel (b)).
@@
According to binary evolution models of spider systems, slower spin periods may result from a lower accretion efficiency during the recycling phase or from a higher initial neutron star mass (e.g., T. M. Tauris et al., 2012; H.-L. Chen et al., 2013; A. Kar et al., 2024; D. Misra et al., 2025b, and references therein). These long-$P$ systems would also exhibit lower spin-down luminosities ($\dot{E}\propto P^{-3}$), leading to reduced companion irradiation. Weaker irradiation inhibits mass loss from the binary, thereby limiting orbital widening. As a result, Cluster 2 spiders would maintain a narrow range of orbital periods while showing a wider distribution of spin periods, shaped by variations in initial neutron star mass or accretion efficiency. In contrast, spiders in Cluster 1 likely underwent more efficient accretion during recycling phase, reaching spin periods of $\sim$2 ms. These systems display higher spin-down luminosities and likely stronger companion irradiation, which lead to wider orbits and a broader range of orbital periods depending on irradiation efficiency.
We note potential outliers for the clustering. Among the RBs, PSR J1302-3258 has $P_{\rm b}=18.8$ hr, $P=3.77$ ms, and $M_{\rm c,min}=0.15M_{\odot}$, standing out in the spin-orbit plane (indicated in Figure 9, panel (b)) and presenting very different derived parameters from Cluster 1 sources ($\dot{E}=5\times 10^{33}$ erg s-1 and $B_{\rm LC}=2.7\times 10^{4}$ G) and shows eclipses at low radio frequencies, fulfilling our RB definition (Section 1.2). However, the eclipses last for only $\sim$10% of the orbit, and the system lacks an optical counterpart despite its apparent proximity ($d\approx 1.4$ kpc from DM and YMW16, Section 3.4), raising doubts about its RB identification (P. Bangale et al., 2024).